sqrt 5x^2-3x=2x can you help me solve this?
Answer by stanbon(75887)

sqrt(5x^2-3x)=2x can you help me solve this?
-----
Square both sides to get:
5x^2-3x = 4x^2
x^2-3x = 0
x(x-3) = 0
----
x = 0 or x = 3
-----
Both answers are good.
Cheers,
Stan H.